OpenStudy (anonymous):

In this question you will derive a general formula for the distance of a point from a line. Let P be the point (p,q) and L the line y=mx+b. The slope of L is . The slope of a line perpendicular to L is . The line through P perpendicular to L can be written as y=sx+c where s is: and c is: . That line intersects L in the point Q=(u,v), where u is: and v is: . The distance of P and Q is . (Note, all answers must be in terms of m, b, p and q.)

OpenStudy (anonymous):

@one2345

OpenStudy (anonymous):

the slope of L is m

OpenStudy (anonymous):

just kidding i didn't get all of it

OpenStudy (anonymous):

The slope of a line perpendicular to L is -1/m?

OpenStudy (anonymous):

the slope of a line that is perpindicular is -1/m

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The line through P perpendicular to L can be written as y=sx+c where s is: ? and c is: ?

OpenStudy (anonymous):

@jayzdd

OpenStudy (anonymous):

s = -1/m

OpenStudy (anonymous):

c= ?

OpenStudy (anonymous):

@jayzdd

OpenStudy (anonymous):

@jayzdd That line intersects L in the point Q=(u,v), where u is: ? and v is: ? The distance of P and Q is ?

OpenStudy (anonymous):

A) m B) -1/m C) line through (p,q) with slope -1/m y - q = -1/m ( x - p) y = -1/m * x + p/m + q s = -1/m c = p/m + q

OpenStudy (anonymous):

whats u and v? @jayzdd

OpenStudy (anonymous):

That line intersects L in the point Q=(u,v), where u is: and v is: . The distance of P and Q is .

OpenStudy (anonymous):

its easier if we graph this 

OpenStudy (anonymous):

so what does that mean

OpenStudy (anonymous):

whats u, v & The distance of P and Q is

OpenStudy (anonymous):

that is what you are given, a point P(x,y) and a line L given by the equation y = mx + b. We want to find the distance from P(x,y) to line L

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

to do this, we draw a perpendicular line through P(x,y) the equation of this perpendicular line is y - q = -1/m ( x - p) solving for y y = -1/m * x + p/m + q

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what is the correct answer for x?

OpenStudy (anonymous):

Now we want to find the intersection of the perp. line and the line L . We will call this intersection point Q(u,v). At the intersection point the y values of the two lines are the same y= mx + b y = -1/m*x + p/m + q set them equal mx + b = -1/m*x + p/m + q solve for x mx +1/m* x = p/m + q - b x ( 1 + 1/m) = p/m + q -b x = ( p/m + q -b) / (1 + 1/m)

OpenStudy (anonymous):

and i still need That line intersects L in the point Q=(u,v), where u is: and v is:

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

whats u & v

OpenStudy (anonymous):

That line intersects L in the point Q=(u,v), where u is: and v is:

OpenStudy (anonymous):

u = ( mq + p - bm) / (1 + m^2)

OpenStudy (anonymous):

whats v?

OpenStudy (anonymous):

@jayzdd

OpenStudy (anonymous):

v = (m^2*q+m*p+b)/(m^2+1)

OpenStudy (anonymous):

u and distance were wrong

OpenStudy (anonymous):

i didnt find the distance yet

OpenStudy (anonymous):

show me a screenshot of the error

OpenStudy (anonymous):

oh what is it

OpenStudy (anonymous):

it just says its not correct for u

OpenStudy (anonymous):

u said u is u = ( mq + p - bm) / (1 + m^2)

OpenStudy (anonymous):

got u

OpenStudy (anonymous):

need distance and then we're done :)

OpenStudy (anonymous):

$$u = \frac{-bm + p + mq} {m^2 + 1}$$

OpenStudy (anonymous):

got u need distance

OpenStudy (anonymous):

its correct now?

OpenStudy (anonymous):

i put it in wrong u had it correct

OpenStudy (anonymous):

i just need the distance now